### v2

# Proto-LeakNet: Towards Signal-Leak Aware Attribution in Synthetic Human Face Imagery

Abstract
The growing sophistication of synthetic image and deepfake generation models has turned source attribution and authenticity verification into a critical challenge for modern computer vision systems. Recent studies suggest that diffusion pipelines unintentionally imprint persistent statistical traces, known as signal-leaks, within their outputs, particularly in latent representations. Building on this observation, we propose Proto-LeakNet, a signal-leak-aware and interpretable attribution framework that integrates closed-set classification with a density-based open-set evaluation on the learned embeddings, enabling analysis of unseen generators without retraining. Acting in the latent domain of diffusion models, our method re-simulates partial forward diffusion to expose residual generator-specific cues. A temporal attention encoder aggregates multi-step latent features, while a feature-weighted prototype head structures the embedding space and enables transparent attribution. Trained solely on closed data and achieving a Macro AUC of 98.13%, Proto-LeakNet learns a latent geometry that remains robust under post-processing, surpassing state-of-the-art methods, and achieves strong separability both between real images and known generators, and between known and unseen ones. The codebase will be available after acceptance.
footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
Figure 1: Proto-LeakNet: given an input image $x$, latent features are extracted from the pretrained Stable Diffusion 2.1 Variational Autoencoder (VAE) in Block 1 (Latent Feature Construction). For each diffusion step $t\!\in\!\{0,5,10\}$, we extract a $z_{t}$ latent which gets normalized with dimension $(4,32,32)$. In Block 2 (Discriminative Backbone), each $z_{t}$ is encoded by a ResNet-18, producing embeddings $\{h_{t}\}$ that are temporally aggregated through the Attention Temporal Pooling module to yield a single representation $\bar{h}$. Block 3 (Prototype-Based Attribution) computes distances between $\bar{h}$ and class prototypes $p_{c,m}$, modulated by a feature-wise gating vector $w$ obtained from a small MLP. The resulting attention-weighted distances are aggregated via a LogSumExp scoring function to produce class probabilities $\pi_{c}$. Symbols: “+” denotes the weighted sum over attention coefficients across timesteps, and “*” indicates the element-wise product between the encoder output and the gating vector.
1 Introduction
The rapid progress of generative models has transformed digital content creation, enabling the synthesis of highly realistic images and videos, also called deepfakes, that are often indistinguishable from authentic ones [goodfellow2014generative, ho2020denoising, rombach2022high]. While these advances have fostered creativity and accessibility, they have also blurred the boundary between real and artificial content, posing serious challenges to media forensics and public trust. As deepfakes proliferate across social, political, and creative domains, distinguishing and attributing their origin has become critical for security, accountability, and digital evidence validation [xie2025deepfake, verdoliva2020media]. Early research in multimedia forensics focused primarily on classifying whether an image is real or generated [guarnera2020deepfake, wang2021fakespotter]. However, a deeper forensic question lies in attribution, namely identifying which generative model produced a given image. Such task, essential for tracing provenance and assessing responsibility, remains extremely challenging, especially in open-set conditions where unknown generators appear at test time [khoo2022deepfake, bindini2024tiny]. Recent studies have shown that even advanced detectors struggle to generalize beyond the closed domain or to maintain interpretability when facing unseen architectures [corvi2023detection, wang2025openset]. Diffusion models have recently equaled most generative models in terms of image quality, yet they introduce subtle statistical artifacts in their latent representations, known as signal-leaks, caused by residual low-frequency information that survives the noising process [Everaert2024SignalLeak]. These traces, although imperceptible, encode model-specific biases and can serve as reliable forensic cues for source attribution. While prior works have improved detection accuracy, they typically lack robustness to domain shifts and offer limited interpretability, especially under open-set or heavily post-processed conditions. To overcome these issues, we aim to exploit the intrinsic statistical biases embedded in diffusion latents as stable, model-specific signatures. In detail, we propose Proto-LeakNet which encodes image $x$ through Stable Diffusion latents and temporal attention, aggregates timestep embeddings via ResNet-18, and performs attribution using prototype-based distances modulated by per-feature attention and gating. The pipeline is synthetically sketched in Fig. 1. Our main contributions are the following:
  • We introduce Proto-LeakNet, a signal-leak-aware and interpretable attribution framework that operates entirely in the latent domain of diffusion models, learning generator-specific biases as stable forensic cues which demonstrates that modeling signal-leak bias in latent space leads to robust attribution
  • We design a temporal attention pooling mechanism that aggregates latent representations across diffusion timesteps, enhancing discriminative power and interpretability by revealing which steps contribute most to attribution.
  • We propose a prototype-based attribution head that shapes the latent geometry through learnable class prototypes and per-feature attention, enabling both compact cluster formation and feature-level interpretability.
  • We develop a density-based open-set evaluation using kernel density estimation on the learned embeddings to assess separability between real images and known generators, and between known and unseen ones without retraining.
This work is organized as follows. Section 2 reviews the relevant state of the art. Section 3 introduces our proposed framework. Section 4 describes the dataset and metrics used in our study. Section 5 reports the experimental results and ablations. Section 6 examines the implications and limitations of our approach. Finally, Section 7 summarizes the key findings and outlines future research directions.
2 Related Work
Beyond binary detection of real vs. fake content [fu2025exploring, cai2025adaptive], recent research has shifted toward identifying the source of generated imagery. Khoo et al. [khoo2022deepfake] outlined the foundations of attribution and categorized deepfake types, identity swap, attribute and expression manipulation, and full-face synthesis, emphasizing interpretability over detection accuracy. Yu et al. [yu2021artificial] proposed a proactive strategy that embeds artificial fingerprints into training data, allowing reliable tracing of model provenance independent of visual artifacts. However, such closed-set settings struggle with unseen generators. Sun et al. [sun2023contrastive] addressed that with the Open-World DeepFake Attribution (OW-DFA) benchmark and the Contrastive Pseudo Learning framework, combining global–local feature voting and confidence-based pseudo-labeling for open-world attribution. Complementary work by Baxevanakis et al. [baxevanakis2025deepfake] showed that standard attribution models generalize poorly across datasets, while contrastive objectives yield more transferable manipulation-aware embeddings. Recent works, for instance, by Tan et al. [freqnet2023] introduce FreqNet, a frequency-domain deepfake detector that forces the model to learn high-frequency features and even operates part of its network in the FFT domain. It includes a dedicated frequency branch between FFT and iFFT to learn source-agnostic cues. Bernabeu-Pérez et al. [susy2024] proposed SuSy, a A CNN-based patch-level forensic classifier that detects whether an image is AI-generated and even attributes it to one of several generator class. It uses a spatial focus: extracting high-contrast patches (via GLCM texture analysis) to train on local artifacts. Wang et al. [latenttracer2024] introduces LatentTracer, a reverse-engineering approach that inverts images into a generative model’s latent space to see if they “belong” to that model. By comparing reconstruction losses, it identifies the source generator without any embedded watermark. Liu et al. [occlip2024] proposes OCC-CLIP, a few-shot one-class attribution framework built upon CLIP embeddings to recognize unseen generators with limited samples. Tan et al. [npr2024] introduces NPR, a method that explicitly targets the telltale upsampling artifacts left by CNN-based generators. It computes local pixel-wise differences (residuals) in small image patches to capture the subtle correlations introduced by upsampling layers. Vasilcoiu et al. [vasilcoiu2025latte] presents LATTE, a transformer-based framework operating directly in the latent space of diffusion models, detecting generated content by modeling long-range dependencies among latent tokens. Finally, Everaert et al. [Everaert2024SignalLeak] identify a signal-leak bias in diffusion latents, showing that low-frequency mismatches encode generator-specific traces. Existing methods still face key limitations, including static latent representations and limited interpretability, often failing to generalize when exposed to unseen generators. Building on these insights, we introduce Proto-LeakNet, a framework that explicitly models and aggregates latent residuals across diffusion steps through prototype supervision and temporal attention. Our approach overcomes these issues by capturing the temporal evolution of generator-specific cues, enabling robust and interpretable attribution under both closed- and open-set conditions.
3 Proposed Method
The proposed Proto-LeakNet framework is deployed into three functional blocks. Latent Feature Construction, extracts informative representations from the diffusion latents of Stable Diffusion 2.1 [rombach2022high]. Discriminative Backbone, aggregates these multi-scale features through a temporal attention encoder. The Prototype-Based Attribution Head interprets the learned representation using class-specific prototypes for interpretable generator attribution. Finally, once defined the pipeline we introduce the generalization methodology.
3.1 Latent Feature Construction
We extract features directly from the latent domain of Stable Diffusion 2.1 (SD2.1). Each image $x\in\mathbb{R}^{3\times H\times W}$ is resized to $256\times 256$ and encoded through the pretrained SD2.1’s Variational Autoencoder (VAE) into a latent $z_{0}$ scaled by the constant $s$ as in
$z_{0}=\frac{\mathcal{E}(x)}{s},$ (1)
where $\mathcal{E}(\cdot)$ is the VAE encoder. To reveal and aggregate residual diffusion traces, we reapply the forward diffusion process at discrete steps $t\in\mathcal{T}=\{0,5,10\}$, sampling
$z_{t}=\alpha_{t}z_{0}+\sigma_{t}\,\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,I),$ (2)
where $t$ denotes the diffusion timestep and $(\alpha_{t},\sigma_{t})$ follow the cosine schedule with
$\alpha_{t}^{2}+\sigma_{t}^{2}=1.$ (3)
Each $z_{t}$ is normalized by $\sigma_{t}$ to maintain consistent scale, producing $\{z_{t}/\sigma_{t}\}_{t\in\mathcal{T}}$. We adopt the early-step set $\{0,5,10\}$ because these timesteps lie in the low-noise regime where generator-specific residuals remain most informative. This choice allows us both to study how the signal leak evolves across noise levels and to temporally aggregate its most discriminative components before they are suppressed by high-variance noise at larger $t$.
Figure 2: Interpretability of Proto-LeakNet (t-SNE Evolution). Layer-wise evolution of the embedding space, illustrating an approximate progression of how prototypes and attention progressively refine and generalize class separation across layers.
3.2 Discriminative Backbone
The feature tensors $\{z_{t}\}_{t\in\mathcal{T}}$ are encoded by a ResNet18 [he2015deepresidual] backbone $\phi(\cdot;\theta)$, with the first convolution adapted to match the channel dimensionality of $z_{t}$, where $\theta$ denotes the trainable backbone parameters. For each timestep $t$, the backbone produces a latent embedding
$h_{t}=\phi(z_{t};\,\theta)\in\mathbb{R}^{D},$ (4)
where $D$ is the embedding dimensionality. A learnable attention module assigns relevance to each timestep via
$a_{t}=\frac{\exp(q^{\top}u_{t})}{\sum_{t^{\prime}\in\mathcal{T}}\exp(q^{\top}u_{t^{\prime}})},\qquad u_{t}=W_{a}h_{t}+b_{a},$ (5)
where $q$, $W_{a}$, and $b_{a}$ are learned parameters. The temporally aggregated embedding is obtained as
$\bar{h}=\sum_{t\in\mathcal{T}}a_{t}\,h_{t},$ (6)
where $\sum_{t}a_{t}=1$. The weights $\{a_{t}\}$ provide temporal interpretability by quantifying the contribution of each diffusion step to the final embedding $\bar{h}$.
3.3 Prototype-Based Attribution Head
Each class $c\!\in\!\{1,\dots,C\}$ is represented by $M$ learnable prototypes, $p_{c,m}\!\in\!\mathbb{R}^{D}$, where $m\in M$, which serve as representative points in latent space. Empirically, four prototypes yielded compact yet well-separated latent clusters. An attention feature-wise gating vector $w\!\in\!(0,1)^{D}$ is computed using a small MLP:
$w=\sigma(A\bar{h}+b),$ (7)
where $A$ and $b$ are learnable parameters. The attention-weighted distance between $\bar{h}$ and each prototype $p_{c,m}$ is defined as
$d_{c,m}(\bar{h},w)=\sum_{i=1}^{D}w_{i}\,(\bar{h}_{i}-p_{c,m,i})^{2}.$ (8)
Per-class scores aggregate distances using a temperature-controlled LogSumExp:
$s_{c}=-\tau\log\sum_{m=1}^{M}\exp\!\Big(-\tfrac{1}{\tau}d_{c,m}(\bar{h},w)\Big),$ (9)
where $\tau>0$ is a learnable scalar that determines the aggregation smoothness. Posterior probabilities are computed as
$\pi_{c}=\frac{\exp(s_{c})}{\sum_{c^{\prime}}\exp(s_{c^{\prime}})},$ (10)
where $c^{\prime}$ denotes the index iterating over all classes in the denominator. The model is trained via cross-entropy loss
$\mathcal{L}_{\text{CE}}=-\frac{1}{B}\sum_{b=1}^{B}\log\pi_{y^{(b)}},$ (11)
where $B$ is the mini-batch size and $y^{(b)}$ are ground-truth labels. All parameters $\{\theta,A,b,p_{c,m}\}$ are optimized jointly with AdamW and weight decay. Supplementary Material will further detail such parameters.
Mahalanobis Scoring and Interpretability
During evaluation, embeddings are scored using a diagonal Mahalanobis comparator fitted on training embeddings. For each class $c$, with empirical mean $\mu_{c}$ and diagonal covariance $\Sigma_{c}=\mathrm{diag}(\sigma^{2}_{c,1},\dots,\sigma^{2}_{c,D})$, the score is
$s_{c}^{\text{maha}}(\bar{h})=-\sum_{i=1}^{D}\frac{(\bar{h}_{i}-\mu_{c,i})^{2}}{\sigma_{c,i}^{2}+\epsilon},$ (12)
where $\sigma_{c,i}$ represents the empirical standard deviation of feature $i$ within class $c$, capturing the intra-class variance along each embedding dimension and $\epsilon$ prevents numerical instability. These scores provide calibrated likelihoods for attribution and open-set evaluation.
Interpretability of Proto-LeakNet
Proto-LeakNet is inherently interpretable. Each distance in Eq. 8 decomposes feature-wise as
$d_{c,m}(\bar{h},w)=\sum_{i=1}^{D}\underbrace{w_{i}(\bar{h}_{i}-p_{c,m,i})^{2}}_{r_{c,m,i}},$ (13)
where $r_{c,m,i}$ measures the contribution of feature $i$. Prototype responsibilities are obtained as
$\pi_{c,m}=\frac{\exp(-d_{c,m}/\tau)}{\sum_{m^{\prime}}\exp(-d_{c,m^{\prime}}/\tau)}.$ (14)
The most activated prototype $\arg\max_{m}\pi_{c,m}$ identifies the latent region that best matches the input. Together, the feature gates $w$, temporal weights $\{a_{t}\}$, and prototype responsibilities $\{\pi_{c,m}\}$ provide a three-level interpretability hierarchy visible in Fig. 2 in detail with the effective distribution.
3.4 Representation-Level Generalization via Density Estimation
While Proto-LeakNet is trained only on closed-set generators, its latent encoder learns a structured representation that can be evaluated for generalization without requiring explicit supervision on unseen classes. In this setting, the goal is not to classify open samples, but to assess whether the learned latent geometry consistently separates embeddings of known generators from those of unseen ones. After training, we discard the prototype-based classifier and use only the frozen ResNet18 backbone-based encoder to produce pooled embeddings $\bar{h}\!\in\!\mathbb{R}^{D}$ for both closed and open samples, forming the sets $\mathcal{H}_{\mathrm{c}}=\{\bar{h}^{(\mathrm{c})}_{i}\}_{i=1}^{N_{\mathrm{c}}}$ and $\mathcal{H}_{\mathrm{o}}=\{\bar{h}^{(\mathrm{o})}_{j}\}_{j=1}^{N_{\mathrm{o}}}$. A Gaussian kernel density estimator (KDE) is fitted on $\mathcal{H}_{\mathrm{c}}$ to model the manifold of closed embeddings:
$p_{\text{KDE}}(h)=\frac{1}{N_{\mathrm{c}}(2\pi\sigma^{2})^{\frac{D}{2}}}\sum_{i=1}^{N_{\mathrm{c}}}\exp\!\left(-\frac{\|h-\bar{h}^{(\mathrm{c})}_{i}\|^{2}}{2\sigma^{2}}\right)$
where $\sigma$ is the kernel bandwidth and $D$ is the embedding dimensionality. For each sample $h$, we compute its log-likelihood score
$s(h)=\log p_{\text{KDE}}(h),$ (16)
which measures how likely $h$ lies within the distribution of known generators. High $s(h)$ values correspond to familiar latent regions, while low scores indicate that the sample is far from any known manifold, suggesting an unseen generator. This approach defines a form of representation-level generalization: the model is never trained on open samples or labels, yet its latent space forms structured low-density regions that naturally reject out-of-distribution inputs. Rather than predicting unseen categories, the KDE analysis evaluates whether the learned representation preserves discriminative geometry under domain shifts. Open generators, though unlabeled, consistently occupy regions separated from the closed manifold, showing that the encoder captures signal-leak statistics that generalize beyond training sources and enable unsupervised detection of unseen generators.
4 Dataset and Evaluation Metrics
Our evaluation builds on WILD [bongini2025wild], a 20k-image benchmark comprising high-quality samples from a closed set of text-to-image generators and an open set of additional state-of-the-art models, with no prompt or model overlap. WILD is selected for its realistic, artifact-free images, preventing attribution bias rooted in visible flaws. The dataset also includes post-processed variants introducing realistic degradations.
Closed set (10k images). The closed set includes ten text-to-image models: Adobe Firefly [firefly], DALL·E 3 [dalle3], FLUX.1 [flux2024], FLUX 1.1 Pro [fluxpro], Freepik [freepik], Leonardo AI [leonardoai], Midjourney [midjourney], Stable Diffusion 3.5 Large [stable_diffusion_35_large], Stable Diffusion XL Turbo [sdxl_turbo], and StarryAI [starryai]. A pool of 1,000 prompts is applied uniformly across all generators. The official split (also used for our experiments) contains 5,000 training images, 2,000 validation images, and 3,000 test images, with splits defined at the prompt level.
Post-Processed Closed-Set. The closed-set post-processed test images underwent 1 (Step 1), 2 (Step 2), and 3 (Step 3) random transformations selected from compression, cropping, resizing, rotation, blur, photometric changes, grayscale conversion, and super-resolution.
Open set (10k images, test-only). The open set includes ten additional generators, again with 1,000 images each: DALL-E, 1 [ramesh2021zeroshot], DeepAI [deepai], HotpotAI [hotpotai], NVIDIA Sana [xie2024sana], Stable Cascade [pernias2023wuerstchen], Stable Diffusion Attend&Excite [sd_ae], StyleGAN [karras2019stylebasedgeneratorarchitecturegenerative], StyleGAN2 [karras2020analyzingimprovingimagequality], StyleGAN3 [karras2021alias], and Tencent Hunyuan [li2024hunyuandit]. This set includes GAN models and text-to-image diffusion/transformer models not found in the closed set. All open-set images were used exclusively for testing.
4.1 Evaluation Metrics
We report two complementary metrics to evaluate both closed- and open-set performance: Macro AUC, Equal Error Rate (EER) and Overlap Coefficient. These jointly measure ranking consistency and separability of confidence scores.
  • Macro AUC: For closed- and open-set scoring, we compute the per-class area under the ROC curve (AUC) (Eq. 17) and average over all $C$ classes:
    $\text{MacroAUC}=\frac{1}{C}\sum_{c=1}^{C}\text{AUC}_{c},$ (17)
    where $\text{AUC}_{c}$ measures the ranking quality for class $c$. This metric evaluates discriminative consistency across classes, independent of the decision threshold.
  • Equal Error Rate (EER): In open-set evaluation, EER is the point at which the false-acceptance rate (FAR) equals the false-rejection rate (FRR), defined in Eq. 18:
    $\text{EER}=\min_{\delta}\big|\text{FAR}(\delta)-\text{FRR}(\delta)\big|,$ (18)
    where $\delta$ denotes the decision threshold. Lower EER indicates better separation between known and unknown samples.
  • Overlap Coefficient (OVL): As shown in Eq. 19, quantifies the empirical intersection between the KDE score distributions of closed and open samples:
    $\text{OVL}=\int\min\big(P_{\text{closed}}(s),\,P_{\text{open}}(s)\big)\,ds,$ (19)
    where $P_{\text{closed}}$ and $P_{\text{open}}$ are normalized density estimates of the log-scores $s$. Values close to zero indicate non-overlapping, perfectly separable distributions.
Table 1: AUC (%) results for each generator class on the raw Closed Set. The table reports both per-class AUC and overall Macro AUC, results in bold are the best, underlined ones are second best.
[…]
5 Experimental Results
We evaluate Proto-LeakNet across multiple scenarios to assess its attribution accuracy, robustness to post-processing, and generalization to unseen data. Our method gets compared against the three best models in WILD [bongini2025wild], FreqNet [freqnet2023], SuSy [susy2024], LatentTracer [latenttracer2024], OCC-CLIP [occlip2024], NPR [npr2024] and LATTE [vasilcoiu2025latte]. Given the similarity in scope and operating domain, we adapted LATTE for source attribution by extending its binary detection objective to a multi-class setting. Owing to its latent-space formulation, the model is naturally predisposed to capture generator-specific cues, making it a suitable baseline for origin attribution. Finally, we include ablation studies to assess the contribution of each component in our pipeline.
5.1 Results on Closed Set
Proto-LeakNet was first evaluated under the closed-set configuration to verify its ability to learn and capture generator-specific signal-leak patterns. All methods to which we compare were trained in our exact conditions. We first analyze the per-class AUCs reported in Table 1. Although Proto-LeakNet does not achieve the top score for every generator, it consistently ranks among the best across all classes, leading to the highest overall Macro AUC. This indicates stronger balance and generalization in capturing generator-specific signal-leak patterns under the closed-set configuration. Compared to all baselines, Proto-LeakNet attains superior latent discriminability and structured separation. The joint effect of temporal attention and prototype supervision is crucial: attention focuses on the most informative diffusion timesteps, while prototypes enforce geometric consistency by attracting samples toward class centroids and repelling those of different generators. Fig. 3 further shows the Top-1 accuracy distributions, where SuSy [susy2024] slightly surpasses our model on raw images (83.32% vs. 82.60%), yet Proto-LeakNet, due to its highest Macro AUC, yields a more compact and interpretable latent manifold as supported by Fig. 2, effectively encoding generator-specific signal-leak cues that support robust and clear attribution.
Table 2: AUC (%) results for each generator class under increasing post-processing levels (Steps 1–3). The last column reports Macro AUC for all models. Results in bold are the best, underlined ones are second best.
[…]
Table 3: Open-set evaluation under different attention configurations. We report AUROC, Equal Error Rate (EER), and Overlap Coefficient (OVL). Lower EER and OVL indicate better separation between closed and open domains.
Configuration AUC (%) EER OVL Attentions Off 57.24 0.44 0.89 Attentions On 56.62 0.45 0.90
5.2 Results on Post-Processed Closed Set
To assess robustness against real-world degradations, all models and Proto-LeakNet were evaluated on progressively post-processed closed-set samples using pretrained checkpoints from the raw configuration. During this test perturbations from Steps 1-3 (Section 4) were introduced only at inference time, without any fine-tuning. Table 2 reports the per-class AUCs, showing that although Proto-LeakNet does not lead every class, it maintains consistently high performance across all generators compared to the other methods, resulting in the best overall Macro AUC at every degradation level. This confirms stronger resilience to visual distortions compared to image-domain baselines, whose performance progressively collapses with increased perturbation. The robustness of Proto-LeakNet stems from its reliance on latent-domain signal-leak cues, which persist even when pixel-level textures are heavily degraded. Fig. 3 further illustrates the Top-1 accuracies across degradation levels. Methods who are based on latent analysis such as OCC-CLIP, LatentTracer, LATTE and Proto-LeakNet obtain the higher metrics compared to pixel-based methods. In detail, Proto-LeakNet achieves the best results overall, being, 75.89%, 74.51%, and 69.52% for Steps 1, 2, and 3 respectively, while the second best method (LATTE), scores 75.64%, 74.27% and 68.60%. In general, these results demonstrate that Proto-LeakNet preserves generator attribution under aggressive post-processing by capturing stable latent biases rather than superficial image artifacts.
Figure 3: Histogram of the Top-1 accuracy distributions per class for each method, from raw to step 1-3.
5.3 Representation-Level Generalization Analysis
To assess how well the learned latent space separates known and unseen data, we evaluate the frozen embeddings of Proto-LeakNet’s ResNet18 encoder rather than its classifier head. This isolates the representational geometry induced by signal-leak cues, independent of class supervision. A kernel density estimator (Section 3.4) is fitted on the closed embeddings and used to measure how open-set unseen generators (Section 4) align with, or deviate from, the learned manifold under three attention configurations. As shown in Fig. 4 (a) and Table 3, in the configuration without attention applied in both embeddings the encoder loses its ability to emphasize generator-specific latent dimensions, resulting in mixed distributions. When attention is applied to both domains as illustrated in Fig. 4 (b), the same weighting pattern is projected onto unseen data, falsely aligning them with the closed distribution and reducing separability. The asymmetric configuration (Fig. 4 (c)), where attention is active only for closed embeddings, preserves generator-specific latent biases and produces distinct, non-overlapping densities indicative of true representation-level generalization. To further validate these observations, we employed the CELEBA-HQ dataset [liu2015faceattributes, karras2018progressive], which contains 30,000 high-quality real face images. With the previous experimental configuration (attention applied only on closed embedding), we treated CELEBA-HQ samples as the open-set domain to test whether the learned embeddings would overlap with the closed generators. The resulting latent distributions (Fig. 4 (d)) remained well separated from the closed clusters, confirming that the learned embedding space effectively discriminates unseen, real data without retraining.
Figure 4: Impact of attention configurations on latent-space separation. (a) Disabling attention for both domains leads to overlapping closed and open clusters. (b) Enabling attention for both aligns open embeddings with the closed distribution. (c) The asymmetric setup, attention active only for closed samples, produces well-separated clusters. (d) Applying the same setup as (c) with real CELEBA-HQ images, separability is preserved.
5.4 Ablation Study
To evaluate the contribution of each component in Proto-LeakNet, we conducted three ablation studies targeting its key modules. First, we examined the role of prototypes and feature-attention in structuring the latent space (Table 4). Removing prototypes substantially disrupts latent organization, lowering both Top-1 Accuracy and Macro AUC, while omitting attention reduces focus on informative latent dimensions. Their combination yields the best results, confirming that prototypes and attention jointly enforce structured, discriminative embeddings that capture signal-leak cues. Replacing Stable Diffusion 2.1 with Stable Diffusion XL (SDXL) produced nearly identical results (Top-1 Acc. 82.55% and Macro AUC of 98.09%), confirming that signal-leak behavior is intrinsic to the diffusion process itself.
Next, we assessed the impact of the encoder architecture (Table 5). Replacing ResNet18 with larger backbones such as EfficientNet-B4, ViT-B16, ResNet50, or ResNet101 decreases performance, indicating that deeper or transformer-based models overfit local latent variations and weaken generalization. ResNet18 provides the best trade-off between compactness and discriminative power, ensuring stable feature extraction across degradations.
We then varied the number of prototypes per class $M$ (Table 6). Using only two prototypes limits intra-class flexibility, while six introduces redundancy and overlap; $M=4$ offers the best balance, yielding compact, well-separated clusters and the highest closed-set performance. Temporal attention pooling, responsible for aggregating diffusion-step features, was not ablated since removing it would collapse temporal aggregation entirely. Moreover, these results validate the design of Proto-LeakNet and highlight how each component contributes to robust and interpretable latent-space attribution.
Table 4: Ablation study on Proto-LeakNet components. We report Top-1 Accuracy and Macro AUC to quantify the effect of removing prototypes and attention mechanisms.
Experiment Top-1 Acc (%) Macro AUC (%) No prototypes 72.63 93.09 No attention 81.23 97.49 No prototypes & no attention 79.80 96.67 Full Proto-LeakNet 82.60 98.13
Table 5: Ablation on different backbone architectures. We report Top-1 Accuracy and Macro AUC on the closed-set configuration to evaluate the impact of the feature extractor on attribution performance.
Backbone Top-1 Acc. (%) Macro AUC (%) EfficientNet_B4 65.80 91.82 ViT-B16 72.07 94.53 ResNet50 81.83 97.39 ResNet101 76.53 95.36 ResNet18 82.60 98.13
Table 6: Ablation on the number of prototypes per class ($M$). We report Top-1 Accuracy and Macro AUC on the closed-set configuration.
Prototypes ($M$) Top-1 Acc. (%) Macro AUC (%) $M=2$ 81.17 97.41 $M=6$ 81.33 97.58 $M=4$ 82.60 98.13
6 Discussion and Limitations
Our study demonstrates that the signal-leak bias constitutes a stable and exploitable forensic cue for source attribution across generative models. Proto-LeakNet represents, to our knowledge, the first framework that explicitly models these latent residuals as discriminative features while providing interpretability across multiple diffusion steps. Across all experiments, the model maintained consistent attribution under post-processing and perturbations, confirming that latent-domain approaches [latenttracer2024, occlip2024, vasilcoiu2025latte] preserve generator-specific cues more effectively than pixel-space methods. Despite its robustness, Proto-LeakNet still presents certain limitations. First, its evaluation is limited to face-centric datasets, leaving open questions about whether signal-leak cues generalize to other visual domains such as objects, scenes, or non-human subjects. Second, the framework relies on the latent representation of Stable Diffusion’s VAE, and it remains unclear whether similar separability would emerge in latents derived from alternative generative families.
7 Conclusions and Future Works
We introduced Proto-LeakNet, a signal-leak-aware attribution framework that combines robustness and interpretability within a latent-space formulation. By operating directly on diffusion latents and modeling residual generator-specific cues, the model learns a structured embedding space shaped by prototypes and attention, enabling reliable attribution even under strong post-processing. Experiments showed that signal-leak bias is a stable and discriminative forensic cue across both diffusion and non-diffusion models, while density-based evaluation revealed clear separability between known and unseen generators. Beyond accuracy, Proto-LeakNet provides transparent prototype- and feature-level explanations linking latent statistics to interpretable evidence. Future work will extend the evaluation to new visual domains and test alternative generative architectures, such as StyleGAN, to assess whether signal-leak traces persist across different encoder families. Treating signal-leak bias as an interpretable fingerprint, establishes a foundation for robust and interpretable attribution of generative media.